# AutoRAN: Automated and Zero-Touch Open RAN Systems

### v3

@@ IV-F Additional Software Components
Another key element of an O-RAN deployment is the RIC, a component that enables observability and control of the network [33]. Specifically, Near-Real-Time (Near-RT) control is achieved via applications called xApps deployed on the Near-RT RIC, while Non-RT control via rApps deployed on the Non-RT RIC We deploy an OSC Near-RT RIC (“E” release) and connect it to OAI via the E2 agent provided in [9, 34] and publicly available as part of the OpenRAN Gym framework [35]. Finally, the cluster also hosts a non-RT RIC to enable observability at larger time scales.
@@ V AutoRAN End-to-End Workflows
In this section, we focus on providing details on how to leverage such blocks to automate the deployment, management and testing of a private 5G network Over-The-Air (OTA). Specifically, we extend our previous work [14] to support a broader set of testing operations and technologies, deploy and test automatically different combinations of protocol stacks and RUs (see Table III) in a repeatable manner, and analyze their performance under different conditions (e.g. number of users, target data rate, and arbitrary protocol configuration files). As we describe later, this is done through a set of pipelines that process high-level, user-specified requirements regarding the specific configuration to test (e.g. OAI with a certain RU and MIMO configuration) and convert them into deployment and testing operations. These involve (i) instantiation of the 5G software containers as pods; (ii) initialization and attachment of UEs; and (iii) data collection of relevant performance metrics to evaluate test success. Additionally, such deployment and testing configurations can be automatically generated through LLMs starting from a high-level intent, and then actuated by the above pipelines, as discussed in Section V-A. To streamline and automate testing procedures and support remote execution of tests, our UEs are Sierra Wireless EM9191 5G modems connected to mini-PCs (e.g., Intel NUC, or Raspberry Pi). The UEs’s mini-PCs are used as host machines that pilot the UE to perform attachment operations via AT commands embedded in the Qualcomm’s chipset, generate traffic and collect data.
@@ V-A Intent-based Instantiation and Testing
  • First, it should be noted that not all the RAN components available in AutoRAN are compatible with each other. For example, OAI supports the ARC L1, while srsRAN does not, meaning we need to prevent the LLM from trying to configure a srsRAN base station with ARC, which would result in a non-working deployment. It is therefore fundamental to ensure that the LLM only selects compatible components and configurations (e.g., OAI with ARC, plain OAI/srsRAN, as shown in Table III). For this reason, we store a compatibility graph, indicating which network components (DU, CU, RU, Core Network) are compatible with one another. Similarly, for each test type, a graph of the list of compatible parameters is fed to the LLM (for example, tests using different traffic generators, such as Multi-Generator (MGEN) can include parameters such as the distribution of traffic, while iPerf tests do not). In the workflow of Figure 7, this graph is consumed when validating values. When a new technology (CU, DU, RU) becomes available and integrated, the dependability graph is updated to reflect the compatibility between that specific element and the others available in AutoRAN.
@@ V-B Deployment Workflow
Each pipeline consists of multiple tasks that need to be executed to successfully run the pipeline. These tasks are associated to operations that span from sending the intent of the user to the LLM, to converting the output of the latter into a deployment object specialized to the physical infrastructure, to configuring and instantiating the gNB components (e.g., CUs and DUs) and connect them to the radio devices used for the test (e.g., either USRP SDRs or commercial RUs) and to the core network. After instantiating the specified components in the form of virtualized containers, the gNB software is started via a call to Flask APIs—or multiple calls in the case of gNB accelerated by the NVIDIA ARC solution, i.e., one for the CU/DU-high container and one for the DU-low container with CUDA Baseband (cuBB)—that they expose. Thus, by leveraging OpenShift’s built-in virtualization and parallelization capabilities, AutoRAN can execute multiple experiments concurrently, provided sufficient hardware resources are available and there is no interference between the tested gNBs.
@@ VI-C End-to-End Deployment and Testing Workflows
In this section, we show experimental results to validate the automated E2E test workflows described in Section V, where we use Sierra Wireless 5G modems as UEs. We leave the extension of the automated tests to smartphone UEs for future work, while we use these devices for manual and mobile tests.
Figure 10 shows the time required to execute the different tasks of the automated testing pipeline discussed in Section V.
@@
Together, the following five bars show the average time required to perform a complete test with five different gNB combinations: (i) GH200 server with ARC; (ii) Gigabyte E251 server with ARC; (iii) Gigabyte E251 server without ARC; (iv) Microway server with OAI and USRP; and (v) Microway server with srsRAN and USRP. The Test Specification bar shows the time taken from when the user specifies which deployment and test configuration to use, to when the gNB Flask API is called to start the test. The pull time, included in this value, is dependent on the size of the image and the network speed. Having a local and private registry inside OpenShift allows AutoRAN to pull the image at the maximum speed allowed by the connection between the cluster and the NAS, that is up to $10$ Gbps. Extra time is still required to the various layers of the image in the target node. On the other hand, after the image has been pulled at least once on a node, it is thereafter kept in its cache registry, and subsequent pulls are in the order of a few milliseconds. Therefore, by pre-loading the image in the internal cache on the node, the pull time becomes negligible even for the bigger images, such as cuBB (around $40$ GB). After the image has been deployed in a pod on the targeted node, the gNB software needs to be instantiated (gNB setup in Figure 10). As shown, the time to run the gNB on the GH200 and Gigabyte servers is quite similar, at around $18$ s. This time accounts for the time to initialize libraries, GPU and NIC for the L1 acceleration. On the other hand, since to run OAI without ARC on the Gigabyte server only one container is needed and the gNB is monolithic, the start time is shorter, about $8$ s. As the RUs is always in idle state (i.e., in the on state but without actively transmitting or receiving) and initialized outside of OpenShift with via M-plane-like functionalities before the gNB pod deployment (see Section IV), we do not include initialization (approximately $60$ s for a Foxconn RU) and reconfiguration times in the instantiation time of CU/DU. This is different from the USRPs deployment, where initialization times are negligible. Indeed, the time taken to start the gNB and connect it to the USRP for both srsRAN and OAI is around $10$ s. Finally, the last two bars show the time taken for the UE (Sierra Wireless 5G modem in our case) to connect to the gNB, and the time taken to collect the results and send them to the data collector pod respectively. It is worth mentioning that we do not report the duration of the data transmissions between gNB and UE, as it depends on the value in the test specifications (e.g., run an iPerf test for $60$ s, see Listing 5). Finally, multiple tests can be executed either as independent tests (e.g., one for iPerf with TCP and one for iPerf3 in UDP) where the gNB is restarted prior to the execution of the next test, or consecutively without redeploying the pod.
@@ VI-D2 Performance Comparison of NVIDIA ARC
In this section, we focus on comparing, for the first time, an OpenShift-based deployment of NVIDIA ARC with its counterpart based on bare-metal and Docker virtualization. In the following, the performance we report is obtained from the cuBB L1 accelerator logs (and not from the UE), since we are interested in comparing the raw throughput being pushed by the accelerator, also considering the partitioned NIC and the extra levels of virtualization. We run a comparison using OTA transmissions with a smartphone and the RuSIM emulator. Results of both experiments are shown in Figure 14. For the former configuration, we generate downlink traffic towards a single Samsung S23 phone UE in a OTA transmission using a Foxconn RU in the N78 band configured at $100$ MHz bandwidth with DDDDDDSUUU as Time Division Duplexing (TDD) pattern. We test iperf3 connectivity between the phone and a server pushing $1000$ Mbps to the phone itself for a duration of $60$ s. The Cumulative Distribution Function (CDF) is calculated as average of 5 experiments. Note that, compared to the Sierra Wireless boards used in the previous tests, this device supports 4 layers in downlink, resulting in higher achievable throughput. The results, indeed, show that the throughput in downlink reaches up to $820$ Mbps, with similar performance for the AutoRAN and bare metal configuration.
To exclude any variability related to the radio environment, we also perform an emulated comparison using RuSIM, one of the Keysight tools available in the Northeastern University OTIC [36]. RuSIM is a RU emulator for 7.2 split, mirroring all parameters that a real RU should expose. In addition to this, it allows simulation of users and of wireless channel. It is usually used jointly with other Keysight tools, including CoreSIM (to emulate core network and traffic generation) and Air Mosaic for managing and orchestrating tests. AutoRAN provides a flexible testing platform and the integration of RuSIM only requires the addition of a second SR-IOV interface to communicate with CoreSIM, since it is external to the cluster. This shows the flexibility of the AutoRAN virtualization capability, since the same physical interface is used twice, for fronthaul and backhaul, using two different virtual functions. We test performance of 15 simulated users, which are able to achieve an aggregate performance of around $1600$ Mbps using all slots ($1600$) every second available over $100$ MHz with a DDDDDDDSUU TDD pattern.
@@
Our results show that the added virtualization and integration into AutoRAN does not lead to relevant performance difference in raw throughput when deploying NVIDIA ARC gNBs. It is worth mentioning that the cuBB logs show some delayed slots from time to time in the OpenShift-based implementation (probably due to the missing optimized kernel), but this situation does not seem to affect performance. Therefore, the AutoRAN approach shows how it is possible to apply infrastructure sharing to the RAN world as in the cloud computing one without significant performance degradations. As at the time of this writing, open-source RAN stacks are not optimized for Ultra Reliable and Low Latency Communications (URLLC) use cases, we evaluate latency by exchanging ICMP packets between a Samsung Galaxy S23 Commercial Off-the-Shelf (COTS) UE and core network on AutoRAN as well as on the bare-metal NVIDIA ARC deployment. Results for this evaluation are shown in Figure 15. We notice that the performance between the original NVIDIA ARC and the AutoRAN deployments are similar, with small variations in latency related to non-deterministic channel conditions and an average ICMP Round Trip Time (RTT) of $18$ ms.
@@ VI-E Coexistence with Different Workloads
This shows how AutoRAN does not introduce relevant delays or performance degradation into the gNB stack while using an additional layer of virtualization and exposing a new set of features that streamlines and simplifies 5G lifecycle management.
@@ VI-F AutoRAN Resilience
The OpenShift container orchestration platform AutoRAN is based upon—which is, in turn, based on Kubernetes—natively provides high-availability mechanisms to compensate for node failures by “evicting” (i.e., moving) workloads from unresponsive to functioning cluster nodes. As mentioned in Section III, however, RAN workloads typically require specific hardware functionalities to be available on the compute nodes. Therefore, they can only move across nodes where these are available (in practice, across nodes part of the same MCP), while an AutoRAN-orchestrated redeployment is needed across nodes with different MCPs. As shown in Fig. 17, the eviction happens automatically after a timeout that starts when OpenShift detects the node to be unavailable. The duration of the timeout is configurable between 0 s (the pod gets evicted as soon as the node appears unhealthy) and infinite (the pod never gets evicted), with a default of 5 minutes. Further analysis of this parameter is left for future work. Further, due to limitations of the protocol stack implementations we leverage, a new instantiation of the pipeline is required to re-start the gNB once the workload is re-allocated. This triggers an RU reset to accommodate the new DU MAC address (alternatively, the DU MAC can be spoofed, but this does not represent a scalable solution without additional middleware, as discussed in [37]). Note that, during this process, UEs can detect a Radio Link Failure (RLF) and connect to a different cell, or directly reconnect to the new gNB, depending on the duration of the eviction timer.

### v4

@@ IV-F Additional Software Components
Another key element of an O-RAN deployment is the RIC, a component that enables observability and control of the network [polese2022understanding]. Specifically, Near-Real-Time (Near-RT) control is achieved via applications called xApps deployed on the Near-RT RIC, while Non-RT control via rApps deployed on the Non-RT RIC We deploy an OSC Near-RT RIC (“E” release) and connect it to OAI via the E2 agent provided in [villa2024x5g, moro2023nfv] and publicly available as part of the OpenRAN Gym framework [bonati2023openrangympawr]. Finally, the cluster also hosts a non-RT RIC to enable observability at larger time scales.
@@ V AutoRAN End-to-End Workflows
In this section, we focus on providing details on how to leverage such blocks to automate the deployment, management and testing of a private 5G network Over-The-Air (OTA). Specifically, we extend our previous work [bonati20235gct] to support a broader set of testing operations and technologies, deploy and test automatically different combinations of protocol stacks and RUs (see Table III) in a repeatable manner, and analyze their performance under different conditions (e.g. number of users, target data rate, and arbitrary protocol configuration files). As we describe later, this is done through a set of pipelines that process high-level, user-specified requirements regarding the specific configuration to test (e.g. OAI with a certain RU and MIMO configuration) and convert them into deployment and testing operations. These involve (i) instantiation of the 5G software containers as pods; (ii) initialization and attachment of UEs; and (iii) data collection of relevant performance metrics to evaluate test success. Additionally, such deployment and testing configurations can be automatically generated through LLMs starting from a high-level intent, and then actuated by the above pipelines, as discussed in Section V-A. To streamline and automate testing procedures and support remote execution of tests, our UEs are Sierra Wireless EM9191 5G modems connected to mini-PCs (e.g., Intel NUC, or Raspberry Pi). The UEs’s mini-PCs are used as host machines that pilot the UE to perform attachment operations via AT commands embedded in the Qualcomm’s chipset, generate traffic and collect data.
@@ V-A Intent-based Instantiation and Testing
  • First, it should be noted that not all the RAN components available in AutoRAN are compatible with each other. For example, OAI supports the ARC L1, while srsRAN does not, meaning we need to prevent the LLM from trying to configure a srsRAN base station with ARC, which would result in a non-working deployment. It is therefore fundamental to ensure that the LLM only selects compatible components and configurations (e.g., OAI with ARC, plain OAI/srsRAN, as shown in Table III). For this reason, we store a compatibility graph, indicating which network components (DU, CU, RU, Core Network) are compatible with one another. Similarly, for each test type, a graph of the list of compatible parameters is fed to the LLM (for example, tests using different traffic generators, such as Multi-Generator (MGEN) can include parameters such as the distribution of traffic, while iPerf tests do not). In the workflow of Figure 7, the compatibility graph is used in the process to validate the configurations generated by the LLM by making sure that those include elements that are compatible with one another. When a new technology (CU, DU, RU) becomes available and integrated, the dependability graph is updated to reflect the compatibility between that specific element and the others available in AutoRAN.
@@ V-B Deployment Workflow
Each pipeline consists of multiple tasks that need to be executed to successfully run the pipeline. These tasks are associated to operations that span from sending the intent of the user to the LLM, to converting the output of the latter into a deployment object specialized to the physical infrastructure, to configuring and instantiating the gNB components (e.g., CUs and DUs) and connecting them to the radio devices used for the test (e.g., either USRP SDRs or commercial RUs) and to the core network. After instantiating the specified components in the form of virtualized containers, the gNB software is started via a call to Flask APIs that they expose. For NVIDIA-ARC deployments, the AutoRAN framework performs a single API call to start the workload. Specifically, an API endpoint is exposed by the CU/DU container. This, then, applies the correct configuration for the deployment and, using an internal connection, starts the DU-low container. Thus, by leveraging OpenShift’s built-in virtualization and parallelization capabilities, AutoRAN can execute multiple experiments concurrently, provided sufficient hardware resources are available and there is no interference between the gNBs.
@@ VI-C End-to-End Deployment and Testing Workflows
In this section, we show experimental results to validate the automated E2E test workflows described in Section V, where we use Sierra Wireless 5G modems as UEs. We leave the extension of the automated tests to smartphone UEs for future work, while we use these devices for manual and mobile tests. Figure 10 shows the time required to execute the different tasks of the automated testing pipeline discussed in Section V.
@@
Together, the following five bars show the average time required to perform a complete test with five different gNB combinations: (i) GH200 server with ARC; (ii) Gigabyte E251 server with ARC; (iii) Gigabyte E251 server without ARC; (iv) Microway server with OAI and USRP; and (v) Microway server with srsRAN and USRP. The Test Specification bar shows the time taken from when the user specifies which deployment and test configuration to use, to when the gNB Flask API is called to start the test. The pull time, included in this value, is dependent on the size of the image and the network speed. Having a local and private registry inside OpenShift allows AutoRAN to pull the image at the maximum speed allowed by the connection between the cluster and the NAS, that is up to $10$ Gbps. Extra time is still required to the various layers of the image in the target node. On the other hand, after the image has been pulled at least once on a node, it is thereafter kept in its cache registry, and subsequent pulls are in the order of a few milliseconds. Therefore, by pre-loading the image in the internal cache on the node, the pull time becomes negligible even for the bigger images, such as cuBB (around $40$ GB). After the image has been deployed in a pod on the targeted node, the gNB software needs to be instantiated (gNB setup in Figure 10). As shown, the time to run the gNB on the GH200 and Gigabyte servers is quite similar, at around $18$ s. This time accounts for the time to initialize libraries, GPU and NIC for the L1 acceleration. On the other hand, since running OAI without ARC on the Gigabyte server only one container is needed and the gNB is monolithic, the start time is shorter, about $8$ s. As the RUs are always in idle state (i.e., in the on state but without actively transmitting or receiving) and initialized outside of OpenShift via M-plane-like functionalities before the gNB pod deployment (see Section IV), we do not include initialization (approximately $60$ s for a Foxconn RU) and reconfiguration times in the instantiation time of CU/DU. This is different from the USRPs deployment, where initialization times are negligible. Indeed, the time taken to start the gNB and connect it to the USRP for both srsRAN and OAI is around $10$ s. Finally, the last two bars show the time taken for the UE (Sierra Wireless 5G modem in our case) to connect to the gNB, and the time taken to collect the results and send them to the data collector pod respectively. It is worth mentioning that we do not report the duration of the data transmissions between gNB and UE, as it depends on the value in the test specifications (e.g., run an iPerf test for $60$ s, see Listing 5). Finally, multiple tests can be executed either as independent tests (e.g., one for iPerf with TCP and one for iPerf3 in UDP) where the gNB is restarted prior to the execution of the next test, or consecutively without redeploying the pod.
@@ VI-D2 Performance Comparison of NVIDIA ARC
In this section, we focus on comparing, for the first time, an OpenShift-based deployment of NVIDIA ARC with its counterpart based on bare-metal and Docker virtualization. In the following, the performance we report is obtained from the cuBB L1 accelerator logs (and not from the UE), since we are interested in comparing the raw throughput being pushed by the accelerator, also considering the partitioned NIC and the extra levels of virtualization. We run a comparison using OTA transmissions with a smartphone and the RuSIM emulator. Results of both experiments are shown in Figure 14. For the former configuration, we generate downlink traffic toward a single Samsung S23 phone UE in an OTA transmission using a Foxconn RU in the N78 band configured at $100$ MHz bandwidth with DDDDDDSUUU as Time Division Duplexing (TDD) pattern. We test iperf3 connectivity between the phone and a server pushing $1$ Gbps to the phone itself for a duration of $60$ s. The Cumulative Distribution Function (CDF) is calculated as average of 5 experiments. Note that, compared to the Sierra Wireless boards used in the previous tests, this device supports 4 layers in downlink, resulting in higher achievable throughput. The results, indeed, show that the throughput in downlink reaches up to $820$ Mbps, with similar performance for the AutoRAN and bare metal configuration.
To exclude any variability related to the radio environment, we also perform an emulated comparison using RuSIM, one of the Keysight tools available in the Northeastern University OTIC [gemmi-otic]. RuSIM is a RU emulator for 7.2 split, mirroring all parameters that a real RU should expose. In addition to this, it allows simulation of users and of wireless channel. It is usually used jointly with other Keysight tools, including CoreSIM (to emulate core network and traffic generation) and Air Mosaic for managing and orchestrating tests. AutoRAN provides a flexible testing platform and the integration of RuSIM only requires the addition of a second SR-IOV interface to communicate with CoreSIM, since it is external to the cluster. This shows the flexibility of the AutoRAN virtualization capability, since the same physical interface is used twice, for fronthaul and backhaul, using two different virtual functions. We test performance of 15 simulated users, which are able to achieve an aggregate performance of around $1600$ Mbps using all slots ($1600$) every second available over $100$ MHz with a DDDDDDDSUU TDD pattern.
@@
Our results show that the added virtualization and integration into AutoRAN does not lead to relevant performance difference in raw throughput when deploying NVIDIA ARC gNBs. It is worth mentioning that the cuBB logs show some delayed slots from time to time in the OpenShift-based implementation (probably due to the missing optimized kernel), but this situation does not seem to affect performance. Therefore, the AutoRAN approach shows how it is possible to apply infrastructure sharing to the RAN world as in the cloud computing one without significant performance degradations.
As at the time of this writing, open-source RAN stacks are not optimized for Ultra Reliable and Low Latency Communications (URLLC) use cases, we evaluate latency by exchanging ICMP packets between a Samsung Galaxy S23 Commercial Off-the-Shelf (COTS) UE and core network on AutoRAN as well as on the bare-metal NVIDIA ARC deployment. Results for this evaluation are shown in Figure 15. We notice that the performance between the original NVIDIA ARC and the AutoRAN deployments are similar, with small variations in latency related to non-deterministic channel conditions and an average ICMP Round Trip Time (RTT) of $18$ ms.
@@ VI-E Coexistence with Different Workloads
In Figure 17, we show the coexistence of multiple RANs on the same node using the SR-IOV and MIG partitioning technologies on the Grace Hopper node. In this experiment, we consider two cells at $100$ MHz: one based on RuSIM, and the other OTA with one COTS UE attached. The L1s of both RANs are completely independent in CPUs, memory, and GPU. We show that the throughput of the first active cell is not impacted by the activation of the second cell, despite sharing the same GPU. Latency is not affected by the coexistence of multiple gNBs, with the same average results as shown in Figure 15. This set of experiments shows how AutoRAN does not introduce relevant delays or performance degradation into the gNB stack while using an additional layer of virtualization and exposing a new set of features that streamlines and simplifies 5G lifecycle management.
Figure 17: Performance comparison with concurrent RAN workloads using NVIDIA ARC.
@@ VI-F AutoRAN Resilience
The OpenShift container orchestration platform AutoRAN is based upon—which is, in turn, based on Kubernetes—natively provides high-availability mechanisms to compensate for node failures by “evicting” (i.e., moving) workloads from unresponsive to functioning cluster nodes. As mentioned in Section III, however, RAN workloads typically require specific hardware functionalities to be available on the compute nodes. Therefore, they can only move across nodes where these are available (in practice, across nodes part of the same MCP), while an AutoRAN-orchestrated redeployment is needed across nodes with different MCPs. The eviction happens automatically after a timeout that starts when OpenShift detects the node to be unavailable. The duration of the timeout is configurable between 0 s (the pod gets evicted as soon as the node appears unhealthy) and infinite (the pod never gets evicted), with a default of 5 minutes. Further analysis of this parameter is left for future work. Additionally, due to limitations of the protocol stack implementations we leverage, a new instantiation of the pipeline is required to re-start the gNB once the workload is re-allocated.
@@
Figure 18 shows an experiment where the OpenShift scheduler detects a failed node (Microway) in around $40$ s. Once this happens, OpenShift enforces a timeout (set at $30$ s), after which the pod is up again and ready to start a new transmission or deployment in a new node of the same MCP. The UE, generating around $30$ Mbps of traffic, gets disconnected, but it reconnects and is able to generate traffic again around $30$ s after the gNB has been redeployed, similarly to what described in Figure 10. Note that, during this process, UEs can detect a Radio Link Failure (RLF) and connect to a different cell, or directly reconnect to the new gNB, depending on the duration of the eviction timer. In the specific case of a 7.2 gNB, the re-deployment of a DU triggers an RU reset to accommodate the new DU MAC address (alternatively, the DU MAC can be spoofed, although this approach would not be scalable without resorting to additional middleware components, as discussed in [foukas2025ranbooster]).
@@ Acknowledgments
[Uncaptioned image] Stefano Maxenti is a Ph.D. Candidate in Computer Engineering at the Institute for the Wireless Internet of Things (WIoT) at Northeastern University, under Prof. Tommaso Melodia. He received a B.Sc. in Engineering of Computing Systems in 2020 and a M.Sc. in Telecommunication Engineering in 2023 from Politecnico di Milano, Italy. His research is linked with AI applications for wireless communications and orchestration, integration, and automation of O-RAN networks.
[Uncaptioned image] Ravis Shirkhani is a Ph.D. Candidate in Computer Engineering at the Institute for the Wireless Internet of Things (WIoT) at Northeastern University, under Prof. Tommaso Melodia. She received a B.Sc. in Electrical Engineering (Communication Systems and Networks) in 2023 from Sharif University of Technology, Iran. Her research focuses on automation of O-RAN networks, power consumption across O-RAN components, and exploring optimization approaches for network energy efficiency.
[Uncaptioned image] Maxime Elkael is a Research Scientist at the Institute for the Wireless Internet of Things at Northeastern University. He received his Ph.D in Computer Science from Institut Polytechnique De Paris/Telecom SudParis in 2023. His research interest lies at the intersection of optimization theory, artificial intelligence and graph theory applied to next generation wireless networks, especially Open RAN networks.
[Uncaptioned image] Leonardo Bonati is an Associate Research Scientist at the Institute for the Wireless Internet of Things, Northeastern University, Boston, MA. He received a Ph.D. degree in Computer Engineering from Northeastern University in 2022. His main research focuses on softwarized approaches for the Open Radio Access Network (RAN) of the next generation of cellular networks, on O-RAN-managed networks, and on network automation, orchestration, and virtualization. He was awarded the 2024 Mario Gerla Award for Research in Computer Science. Leonardo served as TPC co-chair for the IEEE DTwin 2025 workshop, as co-chair for the track on Testbeds, Experimentation and Datasets for Communications and Networking of IEEE CCNC 2025, and as guest editor of the special issue of Elsevier Computer Networks on Advances in Experimental Wireless Platforms and Systems.
[Uncaptioned image] Salvatore D’Oro is the CTO and co-founder of zTouch Networks, a company focused on the development of zero-touch automation solutions for O-RAN systems. He is also a Research Associate Professor at Northeastern University. He received his Ph.D. degree from the University of Catania and is an area editor of Elsevier Computer Communications journal. He serves on the TPC of IEEE INFOCOM, IEEE CCNC & ICC and IFIP Networking. He is one of the contributors to OpenRAN Gym, the first open-source research platform for AI/ML applications in the Open RAN. His research interests include optimization, AI & network slicing for NextG Open RANs.
[Uncaptioned image] Tommaso Melodia is the William Lincoln Smith Chair Professor with the Department of Electrical and Computer Engineering at Northeastern University in Boston. He is also the Founding Director of the Institute for the Wireless Internet of Things and the Director of Research for the PAWR Project Office. He received his Ph.D. in Electrical and Computer Engineering from the Georgia Institute of Technology in 2007. He is a recipient of the National Science Foundation CAREER award. Prof. Melodia has served as Associate Editor of IEEE Transactions on Wireless Communications, IEEE Transactions on Mobile Computing, Elsevier Computer Networks, among others. He has served as Technical Program Committee Chair for IEEE INFOCOM 2018, General Chair for IEEE SECON 2019, ACM Nanocom 2019, and ACM WUWnet 2014. Prof. Melodia is the Director of Research for the Platforms for Advanced Wireless Research (PAWR) Project Office, a \$100M public-private partnership to establish four city-scale platforms for wireless research to advance the US wireless ecosystem in years to come. Prof. Melodia’s research on modeling, optimization, and experimental evaluation of Internet-of-Things and wireless networked systems has been funded by the National Science Foundation, the Air Force Research Laboratory the Office of Naval Research, DARPA, and the Army Research Laboratory. Prof. Melodia is a Fellow of the IEEE and a Distinguished Member of the ACM.
[Uncaptioned image] Michele Polese is a Research Assistant Professor at the Institute for the Wireless Internet of Things, Northeastern University, Boston, since October 2023. He received his Ph.D. at the Department of Information Engineering of the University of Padova in 2020. He then joined Northeastern University as a research scientist and part-time lecturer in 2020. During his Ph.D., he visited New York University (NYU), AT&T Labs in Bedminster, NJ, and Northeastern University. His research interests are in the analysis and development of protocols and architectures for future generations of cellular networks (5G and beyond), in particular for millimeter-wave and terahertz networks, spectrum sharing and passive/active user coexistence, open RAN development, and the performance evaluation of end-to-end, complex networks. He has contributed to O-RAN technical specifications and submitted responses to multiple FCC and NTIA notice of inquiry and requests for comments, and is a member of the Committee on Radio Frequency Allocations of the American Meteorological Society (2022-2024). He is PI and co-PI in research projects on 6G funded by the NTIA, the O-RAN ALLIANCE, U.S. NSF, OUSD, and MassTech Collaborative, and was awarded with several best paper awards and the 2022 Mario Gerla Award for Research in Computer Science. Michele is serving as TPC co-chair for WNS3 2021-2022, as an Associate Technical Editor for the IEEE Communications Magazine, as a Guest Editor in an IEEE JSAC Special Issue on Open RAN, and has organized the Open 5G Forum in Fall 2021 and the NextGenRAN workshop at Globecom 2022.